6)Solve ∂²z/∂x²-∂²z/∂y²=e^{x-2y}+ cos(2x+3y).


1
Expert's answer
2021-10-01T07:22:01-0400

The dependent variable z depends on independent variables x and y.

 p = x z ¶ ¶ , q= y z ¶ ¶ , r= 2 2 x z ¶ ¶ , s= x y z ¶ ¶ ¶ 2 , t= 2 2 y z ¶ ¶

In this case: q + px = x + y is a PDE of order 1

s + t = x2 is a PDE of order 2

Therefore, from the above equation, we obtain: